Two slits separated by a distance of 0.32mm. A beam of 500nm light strikes the slits producing an interference pattern. Determine the number of maxima observed in the angular interval -30°<x<30°
1
Expert's answer
2019-08-20T10:24:28-0400

For double slits the intensity maximum condition is given by equation


"d\sin\theta=m\lambda"

Thus


"m=\frac{d\sin\theta}{\lambda}=\frac{0.32\times 10^{-3}\sin 30^{\circ}}{500\times 10^{-9}}=320"

Therefore in the angular interval -30°<"\theta" <30° there are 640 maxima.